# Leveraging Importance Sampling to Detach Alignment Modules from Large Language Models

### v3

@@ 1 Introduction
In recent years, the rapid advancement of large language models (LLMs) has led to their widespread adoption across various industries[6, 34, 1]. Efforts to align LLM outputs with domain requirements and human values enhance their utility and content safety[29, 11, 23, 28, 2, 17]. Techniques such as supervised learning[36, 41], preference optimization[29, 23, 11] and reinforcement learning[42, 32, 40] are crucial for achieving model alignment.
According to the scaling laws of LLMs[19], increasing model size typically enhances their performance. However, research indicates that effective domain adaptation and value alignment can be achieved even with smaller models[10]. This difference in size requirements necessitates a balance between utility performance and alignment flexibility with respect to model size[37, 30]. Moreover, training large models for specific domains is resource-intensive and requires the deployment of separate models, increasing resource costs and hindering traffic sharing across domains. Thus, there is an urgent need for more efficient and economical model solutions.
Recent works[18, 25, 7] have introduced methods that fine-tune an adapter module on preference datasets to learn correctional residuals between preferred and non-preferred responses, or supervised and synthetic examples, and then stacked onto the upstream model to achieve corrected alignment. While these approaches effectively decouple alignment from LLMs during training, the correction based on the complete upstream response introduces significant latency for the first token during the inference phase, particularly for long content generation. Additionally, the Aligner model $P({\bm{y}}|{\bm{y}}^{\prime},{\bm{x}})$ [18] introduces a reference response $\bm{y}$$^{\prime}$ for correction, which carries the extra potential risk of out-of-distribution (OOD) inputs, not only for the original question $\bm{x}$}, but also for the reference $\bm{y}$$^{\prime}$, as further discussed in Section 4.
@@ 2.2 Detaching the Alignment Module
Equation 3 resembles the structure of the Residual EBM[7], where the controller functions as an energy-based model performing sequence-level alignment. This approach, however, introduces a first-token delay issue similar to that encountered in the Aligner[18]. In the subsequent sections, we will demonstrate that utilizing an autoregressive language model as the alignment module allows for flexible sequence-level training (see Section 2.3) while facilitating minimal time-delay token-level alignment during inference (see Section 2.4).
@@ 2.3 Sequence-level Training
Given any upper bound $\mathrm{U}$ that $\mathbb{E}_{{\bm{x}}\sim{\mathcal{S}},{\bm{y}}\sim P_{\mathrm{M}}}[\log Q_{\theta}({\bm{y}}|{\bm{x}})]\leq\mathrm{U}$, by applying the Lagrange Multiplier Method, we transforms the constrained optimization problem into an unconstrained form:
@@ 2.4 Token-level Aligning
To address this issue, we propose a token-level decoding strategy that leverages the autoregressive properties of both the Proposal Module and the Residual Aligner to reduce first-word delay. Additionally, at each token, we utilize the characteristics of the linear combination of these modules to perform self-normalizing importance sampling[15]. This approach, which we term Proposing-Aligning-Reducing Sampling, effectively circumvents the need for partition function estimation.
@@ 3 Experimental Setup
Model families. To perform importance sampling within the vocabulary space, it is essential that the Proposal Module shares the same vocabulary as the Residual Aligner. This requirement guides our selection of model families, which include multiple models of varying sizes. Specifically, we choose the LLaMA 3 family [14], with model sizes ranging from 1B to 70B, and the Qwen 2.5 family [39], which includes models from 0.5B to 70B. Both families are recognized for their strong performance as leading open-source LLMs. For our experiments, we designate Llama-3.1-8B and Qwen2.5-14B as the Proposal Modules, representing the largest scales that can be trained on a single machine equipped with 8 A800 GPUs within their respective families. These Proposal Modules are paired with Llama-3.2-3B and Qwen2.5-3B as their corresponding Residual Aligners for the main experiments. Additionally, we explore various sizes of Residual Aligners for ablation studies in Section 5.
Tasks and datasets. We conducted experiments on three representative alignment tasks: instruction following, domain adaptation, and preference optimization. For the instruction following task, we randomly selected approximately 120,000 conversations from UltraChat [8], using the first round of chats for training. We utilized the entire TL;DR Summarization dataset [35] for domain adaptation and the complete Anthropic-HH dataset [4] for preference optimization. The details of our experimental datasets are summarized in Table 1.
@@
Training settings. We start the Proposal Module with the original pre-trained model and first conduct a warm-up phase to learn newly introduced special tokens (OOD tokens) in conversation tasks, such as "<|start_header_id|>", "<|end_header_id|>", "<|eot_id|>", etc., particularly within the Llama and Qwen model families. For supervised learning, we use thousands of examples to fine-tune the Proposal Module, enabling it to effectively generate the end-of-sequence token and appropriately conclude conversations. In the case of preference optimization, we follow the approach from [29] to perform SFT using chosen responses. Prior to training, we sample from the Proposal Module across the entire dataset for supervised learning to create the training set. In contrast, the preference optimization allows us to train directly on preference labels, where the chosen response serves as the target and the rejected one acts as the proposal.
@@
Baselines. In supervised learning, we compare the performance of RAM against the Proposal Module that undergoes SFT. To provide a fair comparison, we also include the Aligner [18] and SFT models of equivalent size to the Residual Aligner as baselines, ensuring comparable computational loads.
@@
Evaluation settings. We evaluate our models using the widely recognized open-ended benchmark, AlpacaEval 2 [9], which assesses conversational capabilities across 805 questions sourced from five datasets. We report scores according to the benchmark’s evaluation protocol, employing Qwen2.5-72B-Instruct and GPT-4-1106-preview as evaluators (referred to as Qwen2.5-Eval and GPT4-Eval). Our evaluation includes both length-controlled win rates (LC), which are designed to mitigate the effects of model verbosity, and raw win rates (WR).
@@ 4 Experimental Results
In contrast, the Aligner method, constrained by the SFT framework, is at risk of overfitting and its inference capabilities rely solely on its fewer parameters, limiting performance, particularly at the 1B scale. Consequently, the Aligner tends to generate repetitive patterns [21] and struggles to effectively capture long-context information [13]. These limitations hinder the overall performance of the Aligner, causing it to consistently fall short of the results achieved by the upstream warmed-up and SFT Proposal Modules, especially within the Llama3 family.
@@
We also conduct supplementary experiments to compare our model with Controlled Decoding (CD) [24], emphasizing both output quality and length in Appendix F.1. Additionally, we assess first-token latency in comparison to Aligner in Appendix F.2.
@@ 6 Related Works
Alignment of LLMs. Aligning LLMs with human values is essential for improving their utility and safety. This process has progressed from prompt engineering [22, 38, 12] to systematic methods like alignment tuning. Key techniques include supervised learning, which uses instruction-response paired datasets for supervised fine-tuning (SFT) [41, 36, 33], and reinforcement learning from human feedback (RLHF) [26, 32, 5], which optimizes models based on user preferences but can be complex and resource-intensive. Direct Preference Optimization (DPO) [29] offers a simpler offline alternative by utilizing preference data directly. Various alignment strategies have emerged from DPO’s modeling of rewards, such as Identity-PO (IPO) [3], which replaces unbounded mapping with identity mapping to reduce overfitting, and SimPO [23], which eliminates the reference model in DPO and introduces a length-control mechanism. We propose a method for transferring supervised learning to our Residual Alignment Model by training a smaller alignment module as a residual complement to the larger model, achieving an efficient and flexible solution for aligning large-scale models.
Residual Correction for LLMs. Residual Energy-Based Models (Residual EBMs) [27, 7] enhance text generation by modeling the energy landscape to improve output coherence and control. They build on Energy-Based Models (EBMs) [16, 20, 31] by integrating globally normalized EBMs with local language models, refining a base distribution through energy-based adjustments to capture missed dependencies. Controlled Decoding (CD) [24] also takes the form of Residual EBMs which solve a KL-regularized RL objective to learn a prefix scorer for the reward that is used to steer the generation from a partially decoded path. The Aligner [18, 25] fine-tunes an adapter module on preference datasets to learn correctional residuals between preferred and non-preferred responses, stacking this onto the upstream model for corrected alignment. While effective in decoupling alignment from LLMs during training, the reliance on complete upstream responses introduces significant latency for the first token during inference. Additionally, the Aligner’s use of a reference response poses risks with out-of-distribution inputs. Our method leverages importance sampling to derive the residual alignment module, directly modeling conditional probabilities along with strategies during training and inference to reduce variance and enhance stability with biased estimators. Additionally, we introduce a token-level decoding strategy to achieve minimal first word latency, enhancing usability in practical applications.
@@ Appendix D Implementation of Training and Inference
During the training and inference processes, we maintain consistency in the sampling parameters for the proposal model with those used for the upstream model in Aligner [18], detailed in Table 5, except for the repetition penalty, which aligns with the sampling parameters employed during the inference stage.
@@ F.1 Comparison to Controlled Decoding (CD)
Controlled Decoding (CD) [24] takes the form of Residual EBMs [27, 7] which solve a KL-regularized RL objective to learn a prefix scorer for the reward that is used to steer the generation from a partially decoded path. The prefix scorer evaluates the scores of any sequence prefix, addressing the limitation of Residual EBMs that require evaluation of the entire sequence, thus enhancing practical applicability.

### v4

@@ 1 Introduction
In recent years, the rapid advancement of large language models (LLMs) has led to their widespread adoption across various industries[bi2024deepseek, team2024gemini, achiam2023gpt]. Efforts to align LLM outputs with domain requirements and human values enhance their utility and content safety[rafailov2024direct, ethayarajh2024kto, meng2024simpo, qi2024safety, askell2021general, huang2025trustworthiness, zhu2025fly, liu2025mitigating]. Techniques such as supervised learning[wang2022self, zhou2024lima, zhu2024flipguard], preference optimization[rafailov2024direct, meng2024simpo, ethayarajh2024kto] and reinforcement learning[ziegler2019fine, stiennon2020learning, zhu2024lire, yang2024rlcd] are crucial for achieving model alignment.
According to the scaling laws of LLMs[kaplan2020scaling], increasing model size typically enhances their performance. However, research indicates that effective domain adaptation and value alignment can be achieved even with smaller models[eldan2023tinystories]. This difference in size requirements necessitates a balance between utility performance and alignment flexibility with respect to model size[wei2022chain, ranaldi2024aligning]. Moreover, training large models for specific domains is resource-intensive and requires the deployment of separate models, increasing resource costs and hindering traffic sharing across domains. Thus, there is an urgent need for more efficient and economical model solutions.
Recent works[ji2024aligner, ngweta2024aligners, deng2020residual] have introduced methods that fine-tune an adapter module on preference datasets to learn correctional residuals between preferred and non-preferred responses, or supervised and synthetic examples, and then stacked onto the upstream model to achieve corrected alignment. While these approaches effectively decouple alignment from LLMs during training, the correction based on the complete upstream response introduces significant latency for the first token during the inference phase, particularly for long content generation. Additionally, the Aligner model $P({\bm{y}}|{\bm{y}}^{\prime},{\bm{x}})$ [ji2024aligner] introduces a reference response $\bm{y}$$^{\prime}$ for correction, which carries the extra potential risk of out-of-distribution (OOD) inputs, not only for the original question $\bm{x}$}, but also for the reference $\bm{y}$$^{\prime}$, as further discussed in Section 4.
@@ 2.2 Detaching the Alignment Module
Equation 3 resembles the structure of the Residual EBM[deng2020residual], where the controller functions as an energy-based model performing sequence-level alignment. This approach, however, introduces a first-token delay issue similar to that encountered in the Aligner[ji2024aligner]. In the subsequent sections, we will demonstrate that utilizing an autoregressive language model as the alignment module allows for flexible sequence-level training (see Section 2.3) while facilitating minimal time-delay token-level alignment during inference (see Section 2.4).
@@ 2.3 Sequence-level Training
Given any upper bound $\mathrm{U}$ that $\mathbb{E}_{{\bm{x}}\sim{\mathcal{S}},{\bm{y}}\sim P_{\mathrm{M}}}[\log Q_{\theta}({\bm{y}}|{\bm{x}})]\leq\mathrm{U}$, by applying the Lagrange Multiplier Method, we transform the constrained optimization problem into an unconstrained form:
@@ 2.4 Token-level Aligning
To address this issue, we propose a token-level decoding strategy that leverages the autoregressive properties of both the Proposal Module and the Residual Aligner to reduce first-word delay. Additionally, at each token, we utilize the characteristics of the linear combination of these modules to perform self-normalizing importance sampling[grover2019bias]. This approach, which we term Proposing-Aligning-Reducing Sampling, effectively circumvents the need for partition function estimation.
@@ 3 Experimental Setup
Model families. To perform importance sampling within the vocabulary space, it is essential that the Proposal Module shares the same vocabulary as the Residual Aligner. This requirement guides our selection of model families, which include multiple models of varying sizes. Specifically, we choose the LLaMA 3 family [grattafiori2024llama], with model sizes ranging from 1B to 70B, and the Qwen 2.5 family [yang2024qwen2], which includes models from 0.5B to 70B. Both families are recognized for their strong performance as leading open-source LLMs. For our experiments, we designate Llama-3.1-8B and Qwen2.5-14B as the Proposal Modules, representing the largest scales that can be trained on a single machine equipped with 8 A800 GPUs within their respective families. These Proposal Modules are paired with Llama-3.2-3B and Qwen2.5-3B as their corresponding Residual Aligners for the main experiments. Additionally, we explore various sizes of Residual Aligners for ablation studies in Section 5.
Tasks and datasets. We conducted experiments on three representative alignment tasks: instruction following, domain adaptation, and preference optimization. For the instruction following task, we randomly selected approximately 120,000 conversations from UltraChat [ding2023enhancing], using the first round of chats for training. We utilized the entire TL;DR Summarization dataset [volske2017tl] for domain adaptation and the complete Anthropic-HH dataset [bai2022training] for preference optimization. The details of our experimental datasets are summarized in Table 1.
@@
Training settings. We start the Proposal Module with the original pre-trained model and first conduct a warm-up phase to learn newly introduced special tokens (OOD tokens) in conversation tasks, such as "<|start_header_id|>", "<|end_header_id|>", "<|eot_id|>", etc., particularly within the Llama and Qwen model families. For supervised learning, we use thousands of examples to fine-tune the Proposal Module, enabling it to effectively generate the end-of-sequence token and appropriately conclude conversations. In the case of preference optimization, we follow the approach from [rafailov2024direct] to perform SFT using chosen responses. Prior to training, we sample from the Proposal Module across the entire dataset for supervised learning to create the training set. In contrast, the preference optimization allows us to train directly on preference labels, where the chosen response serves as the target and the rejected one acts as the proposal.
@@
Baselines. In supervised learning, we compare the performance of RAM against the Proposal Module that undergoes SFT. To provide a fair comparison, we also include the Aligner [ji2024aligner] and SFT models of equivalent size to the Residual Aligner as baselines, ensuring comparable computational loads.
@@
Evaluation settings. We evaluate our models using the widely recognized open-ended benchmark, AlpacaEval 2 [dubois2024length], which assesses conversational capabilities across 805 questions sourced from five datasets. We report scores according to the benchmark’s evaluation protocol, employing Qwen2.5-72B-Instruct and GPT-4-1106-preview as evaluators (referred to as Qwen2.5-Eval and GPT4-Eval). Our evaluation includes both length-controlled win rates (LC), which are designed to mitigate the effects of model verbosity, and raw win rates (WR).
@@ 4 Experimental Results
In contrast, the Aligner method, constrained by the SFT framework, is at risk of overfitting and its inference capabilities rely solely on its fewer parameters, limiting performance, particularly at the 1B scale. Consequently, the Aligner tends to generate repetitive patterns [li2023repetition] and struggles to effectively capture long-context information [godey2024small]. These limitations hinder the overall performance of the Aligner, causing it to consistently fall short of the results achieved by the upstream warmed-up and SFT Proposal Modules, especially within the Llama3 family.
@@
We also conduct supplementary experiments to compare our model with Controlled Decoding (CD) [mudgal2024controlled], emphasizing both output quality and length in Appendix F.1. Additionally, we assess first-token latency in comparison to Aligner in Appendix F.2.
@@ 6 Related Works
Alignment of LLMs. Aligning LLMs with human values is essential for improving their utility and safety. This process has progressed from prompt engineering [lin2023unlocking, xie2021explanation, ganguli2023capacity] to systematic methods like alignment tuning. Key techniques include supervised learning, which uses instruction-response paired datasets for supervised fine-tuning (SFT) [zhou2024lima, wang2022self, sun2024principle], and reinforcement learning from human feedback (RLHF) [ouyang2022training, stiennon2020learning, bai2022constitutional], which optimizes models based on user preferences but can be complex and resource-intensive. Direct Preference Optimization (DPO) [rafailov2024direct] offers a simpler offline alternative by utilizing preference data directly. Various alignment strategies have emerged from DPO’s modeling of rewards, such as Identity-PO (IPO) [azar2024general], which replaces unbounded mapping with identity mapping to reduce overfitting, and SimPO [meng2024simpo], which eliminates the reference model in DPO and introduces a length-control mechanism. We propose a method for transferring supervised learning to our Residual Alignment Model by training a smaller alignment module as a residual complement to the larger model, achieving an efficient and flexible solution for aligning large-scale models.
Residual Correction for LLMs. Residual Energy-Based Models (Residual EBMs) [parshakova2019global, deng2020residual] enhance text generation by modeling the energy landscape to improve output coherence and control. They build on Energy-Based Models (EBMs) [hinton2002training, lecun2006tutorial, ranzato2007unified] by integrating globally normalized EBMs with local language models, refining a base distribution through energy-based adjustments to capture missed dependencies. Controlled Decoding (CD) [mudgal2024controlled] also takes the form of Residual EBMs which solve a KL-regularized RL objective to learn a prefix scorer for the reward that is used to steer the generation from a partially decoded path. The Aligner [ji2024aligner, ngweta2024aligners] fine-tunes an adapter module on preference datasets to learn correctional residuals between preferred and non-preferred responses, stacking this onto the upstream model for corrected alignment. While effective in decoupling alignment from LLMs during training, the reliance on complete upstream responses introduces significant latency for the first token during inference. Additionally, the Aligner’s use of a reference response poses risks with out-of-distribution inputs. Our method leverages importance sampling to derive the residual alignment module, directly modeling conditional probabilities along with strategies during training and inference to reduce variance and enhance stability with biased estimators. Additionally, we introduce a token-level decoding strategy to achieve minimal first word latency, enhancing usability in practical applications.
@@ Appendix D Implementation of Training and Inference
During the training and inference processes, we maintain consistency in the sampling parameters for the proposal model with those used for the upstream model in Aligner [ji2024aligner], detailed in Table 5, except for the repetition penalty, which aligns with the sampling parameters employed during the inference stage.
@@ F.1 Comparison to Controlled Decoding (CD)
Controlled Decoding (CD) [mudgal2024controlled] takes the form of Residual EBMs [parshakova2019global, deng2020residual] which solve a KL-regularized RL objective to learn a prefix scorer for the reward that is used to steer the generation from a partially decoded path. The prefix scorer evaluates the scores of any sequence prefix, addressing the limitation of Residual EBMs that require evaluation of the entire sequence, thus enhancing practical applicability.